Density of a gas is found to be 5.46g/L at 27c and 2 bar pressure what is its density at stp
 

Grade:11

1 Answers

Arun
6 years ago
3 = 3 g/dm3Density at STP = mass/volume = m/1.82V1 = (5.46/1.82) g/dm2 = V11.82V2 = V1/273(273/150)V2/300 = 1V12V2/T2V2 = P1/T1V1P